OpenStudy (anonymous):

Find the indicated derivative: 3y + pisin(x^2 y) = pi at (1, pi/6) Thanks. Find the indicated derivative: 3y + pisin(x^2 y) = pi at (1, pi/6) Thanks. @Mathematics

OpenStudy (anonymous):

\[3y +\pi \sin(x ^{2}y) = \pi \]

myininaya (myininaya):

(3y)'=3(y)'=3y' sin(g(x))=g'(x)cos(g(x)) if g(x)=x^2y then g'(x)=(x^2y)'=(x^2)'y+x^2(y)' = 2xy+x^2y' (constant)'=0

OpenStudy (anonymous):

\[3 + \cos[x^2dy/dx + 2xy] = 0 \] by using the product rule I came up with the top ^ and then I simplified and solved for dy/dx though I keep getting the wrong answer, would I be on the right track though?

OpenStudy (anonymous):

\[3y'+\pi\cos(x^2y)(2xy+x^2y')=0\]and we solve for y', or else plug in the numbers now and solve for y' second

OpenStudy (anonymous):

looks like a mistake in your answer above because you have 3 instead of 3y', and also the pi is missing as a factor of the second term

OpenStudy (anonymous):

probably easiest now to replace x by 1 and y by \[\frac{\pi}{6}\] and see what we get.

OpenStudy (anonymous):

\[3y'+\pi\cos(\frac{\pi}{6})(\frac{\pi}{3}+y')=0\]

OpenStudy (anonymous):

now algebra. hope it works

OpenStudy (anonymous):

Ohhh okay, alright I think I get it now, thanks a lot for your help. I appreciate it !

OpenStudy (anonymous):

yw. i got an answer but it sure looks ugly!

OpenStudy (anonymous):

lol hmm the answer should be \[(\pi ^{2}\sqrt{3}) / (18 + 3 \pi \sqrt{3}) \]

OpenStudy (cwrw238):

this is a long tedious differentiation ...

OpenStudy (anonymous):

Yes, I hate these questions, its going to be all over my midterm :S

OpenStudy (cwrw238):

i got dy/dx = -2 pi yx cox(yx^3) / pi x^2 cos(yx^2) + 3

OpenStudy (anonymous):

ohh any chance you could show me how you did that?

OpenStudy (cwrw238):

sorry its yx^2 not yx^3

OpenStudy (cwrw238):

right - i did it on pencil/paper - i'll take another look and try to explain it - it is long winded

OpenStudy (anonymous):

alright thanks a lot.

OpenStudy (cwrw238):

sorry - i've been called away - be back later

OpenStudy (anonymous):

np, thanks for your help anyways

OpenStudy (cwrw238):

3y + pi sin (x^2y) =pi 3 dy/dx + d(pi sin (x^2y) /dx = 0 find d(pi sin (x^2y) /dx: = (x^2 dy/dx + 2xy) * pi cos(x^2y) - by product / chain rule = pi cos(x^2y) * x^2 dy/dx + pi cos(x^2y) * 2xy so 3 dy/dx + pi cos(x^2y) * x^2 dy/dx + pi cos(x^2y) * 2xy = 0 3 dy/dx + pi cos(x^2y) * x^2 dy/dx = - 2 pi xy cos(x^2y) dy/dx = [ -2pi xy cos (x^2 y)] / [pi x^2 cos(x^2 y) + 3]

OpenStudy (cwrw238):

value at (1.pi/6): = -2 pi (pi/6) cos (pi/6) / pi cos (pi/6) + 3 = -2 (pi)^2/6 * sqrt3/2 ---------------- sqrt3 pi/ 3 + 3 = - sqrt3 (pi)^2 2 ----------- * ------ 6 sqrt3 pi + 6 = -sqrt 3 (pi(^2 ) / ( 3 sqrt3 pi + 18)